Product of two consecutive even integers plus their sum is 98. Find the numbers ?
Found 2 solutions by lenny460, richwmiller:
Answer by lenny460(1073)

Product of two consecutive even integers plus their sum is 98. Find the numbers ?
Let the two numbers be n and n + 2.
Therefore:
n + n + 2 = 98
2n + 2 = 98
2n = 96
n = 48

Answer:
The two numbers are 48 and 50






lennox obuong
algebra tutor

Answer by richwmiller(17219)

x*y+x+y=98,
y=x+2
x^2+4x-96 = 0
(x-8)*(x+12)
x=-12 y=-10
x = 8, y = 10
check
8*10=80
8+10=18
80+18=98
-12*-10=120
-12+-10=-22
120+(-22)=98